OpenStudy (kaylardz0405):

Simplify the expression. 3 • 7 • 23 A. 168 B. 21 C. 84 D. 42

OpenStudy (kaylardz0405):

wanted to double check my answer is it A

Nnesha (nnesha):

simplify what ?

OpenStudy (kaylardz0405):

sorry 3*7*2 to the 3rd power

Nnesha (nnesha):

\[\huge\rm (3 \times 7 \times 23)^3\] like this ?

OpenStudy (kaylardz0405):

no just 2 to the 3rd power not 23

Nnesha (nnesha):

\[\huge\rm (3 \times 7 \times 2)^3\] like this ?

OpenStudy (kaylardz0405):

yes

OpenStudy (kaylardz0405):

is it A

Nnesha (nnesha):

how is it A ??

OpenStudy (kaylardz0405):

2 to the 3rd power is 8

OpenStudy (kaylardz0405):

8*7= 56*3=168

Nnesha (nnesha):

\[\huge\rm 3 \times 7 \times (2)^3\] is it your question right ?

OpenStudy (kaylardz0405):

yes

Nnesha (nnesha):

it's not (3 times 7 times 2)^3 TOTALLY DIFFERENT

Nnesha (nnesha):

yes right it's 168

OpenStudy (kaylardz0405):

ohh then can you expain it to me please

OpenStudy (kaylardz0405):

so i am right ???

Nnesha (nnesha):

\[\huge\rm ( 3 \times 7 \times 2)^3 = (3)^3 \times (7)^3 \times (2)^3\] (3 times 7 times 2) all 3 numbers to the 3rd meaning 3 is to the 3rd power 7 is to the 3rd power and 2 as well

Nnesha (nnesha):

yes you are right

OpenStudy (kaylardz0405):

oh sorry only 2is in parenthisis

Nnesha (nnesha):

yeah you got it!

OpenStudy (kaylardz0405):

ok thanks

Nnesha (nnesha):

np :-)